# Atomistic modeling of uranium monocarbide with a machine learning interatomic potential

### v1

@@ IV Conclusions
We present the first MLIP specifically developed for uranium monocarbide, trained on high-quality DFT+U data and validated across critical material properties. The MLIP demonstrates robust predictive capabilities, accurately reproducing lattice parameters, elastic constants, defect energetics, and diffusion mechanisms across a range of thermodynamic conditions. While some activation energies for defect diffusion are moderately overestimated, relative trends remain consistent with experimental observations, particularly emphasizing the importance of defect-driven transport. This MLIP represents a significant step toward improving realistic atomistic simulations that can lead to a better understanding of UC in reactor conditions and it is further qualification as a fuel.
@@ Acknowledgements.
We calculated the specific heat capacity at constant pressure, shown in Fig. S1 and bulk moduli, shown in Fig. S1 as a function of temperature.
Since MD calculations only capture the vibrational component of the heat capacity–excluding electronic (quantum) and magnetic contributions– there is a large discrepancy with the experimental-inferred heat capacity [1], particularly at low temperatures, where the quantum effects are stronger. However, the MLIP captures the increase in heat capacity at high temperatures, seen in UC, and correlated to the thermally activated point defects [2]. The bulk moduli calculated at zero-pressure, correctly predicts the inverse relation between bulk modulus and lattice expansion. The value at 300 K, agrees with the experimental data reported around 156–158 GPa.
@@
Figure S1: MLIP predicted thermophysical properties as function of temperature. Specific heat capacity with experimental data from Ref. [1] (left), and bulk moduli (right).

### v2

@@ IV Conclusions
We present the first MLIP specifically developed for uranium monocarbide, trained on high-quality DFT+U data and validated across critical material properties. The MLIP demonstrates robust predictive capabilities, accurately reproducing lattice parameters, elastic constants, defect energetics, and diffusion mechanisms across a range of thermodynamic conditions. While some activation energies for defect diffusion are moderately overestimated, relative trends remain consistent with experimental observations, particularly emphasizing the importance of defect-driven transport. This MLIP represents a significant step toward performing realistic atomistic simulations, as it achieves greater accuracy than classical potentials with lower costs than DFT calculations. This MLIP can lead to a better understanding of UC in reactor conditions and help with qualification of UC as an alternative nuclear fuel candidate.
@@ Acknowledgements.
t-sne visualization of clusters
Figure S1: Two-dimensional t-SNE visualization of the UC dataset after defect-based classification identified using a lattice-based analysis: stoichiometric distorted (expanded, mid-range, and compressed), Frenkel pairs, and neutral vacancy configurations. Structures are color coded based on volume per atom.
Figure S1 presents a t-distributed stochastic neighbor embedding (t-SNE) visualization of the dataset, where points are colored based on the volume per atom. The data display a clear clustering between structure and volumetric strain. Atomic configurations are represented using a global structural descriptor based on interatomic distance distributions. Defect-containing configurations, including Frenkel defects and bivacancies, span a broader range of volumes, reflecting the local strain and lattice distortions introduced by point defects.
Temperature-dependent heat capacity and bulk modulus
Fig. S2 presents the specific heat capacity at constant pressure. Fig. S2 presents the bulk moduli as a function of temperature. Both figures compare the HIP-NN MLIP predictions with available experimental data.
Since MD calculations only capture the vibrational component of the heat capacity–excluding electronic (quantum) and magnetic contributions– there is a large discrepancy with the experimental-inferred heat capacity [1]. As depicted in Fig. S2, the MLIP shows a steady increase of the heat capacity with a sudden increase after 2000 K, that is not reflected in the experimental data. The bulk moduli calculated at zero-pressure, correctly predicts the expected inverse relation between bulk modulus and lattice expansion (see Fig. S2). The MLIP bulk modulus at 300 K is lower than the experimental data, which is around 156–158 GPa for stoichoimetric UC [2].
@@
Figure S2: MLIP predicted thermophysical properties as function of temperature. (a) Specific heat capacity with experimental data from Ref. [1], and (b) Bulk moduli with experimental data from Ref. [2].
Potential performance with increased tensor sensitivity
Constructing an MLIP often requires tedious hyperparameter tuning. One key hyperparameter for HIP-NN is the order of tensor sensitivity $({\ell_{\text{max}}})$. Although higher tensor sensitivity often results in lower training errors, this comes at a greater computational cost for running MD and potentially a higher risk of overfitting. Thus, in this section, we investigate the impact of increasing the HIP-NN tensor sensitivity on the model performance, in particular, on the phonon dispersion, elastic constants, and defect energetics. For this purpose, we train a model with quadrupole tensor sensitivity (${\ell_{\text{max}}}=2$, HipnnQuad) and compare with the model presented in the main text that utilizes vector tensor sensitivity (${\ell_{\text{max}}}=1$, HipnnVec). For both models, we set the cusp regularization to 0.01 as this value is essential to ensure mechanical stability. Specifically, a HIP-NN ${\ell_{\text{max}}}=2$ model with a lower cusp regularization ($1e^{-6}$) exhibited unphysical elastic behavior.
Table S1: Elastic constants (GPa) obtained from MLIP models compared to DFT+$U$ and a classical potential.
Model $C_{11}$ $C_{12}$ $C_{44}$
HipnnVec 296.6 84.2 63.5
HipnnQuad 297.8 91.4 28.4
DFT+$U$ 333 73 48
By comparing the elastic constants (Table S1) for both MLIPs, we find that the $C_{11}$ and $C_{12}$ values are quite similar for HipnnVec and HipnnQuad. However, the HipnnQuad model substantially underestimates the shear modulus $C_{44}$, indicating an overly soft response to shear deformations.
The phonon dispersions, shown in Fig. S3, reveal that the quadrupole model (HipnnQuad) reproduces optical phonon branches more closely than the vector model (HipnnVec) along high-symmetry paths. The observed trade-off between optical phonon accuracy and elastic stability underscores that phonon agreement away from $\Gamma$ is not sufficient to guarantee physically correct elastic behavior.
[…]
Table S2 summarizes Schottky defect and Frenkel energies for carbon and uranium predicted by the two MLIPs, compared against DFT+$U$. The HipnnQuad model reproduces defect formation energies more closely than the HipnnVec model, by approximately 0.3 to 0.7 eV. Defect formation energies are dominated by highly localized distortions, which are well captured by higher-order angular sensitivity channels available in the quadrupole model.
Table S2: Point-defect formation energies (eV) predicted by MLIP models compared to reference data.
Model SD $_{\text{C}} FP$$_{\text{U}}$
HipnnVec 5.58 4.41 9.60
HipnnQuad 5.22 4.15 8.84
DFT+$U$ 4.3 3.4 8.1
Based on the combined assessment of elastic constants, phonon stability and defect energetics the HipnnVec model provides a more balanced consistent description of uranium carbide. While HipnnQuad yields improved optical phonon agreement and defect formation energies, HipnnQuad is extremely deficient in predicting the shear response. Our decision to utilize HipnnVec as a general-purpose production potential is motivated by the similar accuracy to HipnnQuad and the lower computational cost for running MD.